OpenStudy (anonymous):

absolute value 7-12/12-7

jimthompson5910 (jim_thompson5910):

\[\left| \frac{7-12}{12-7} \right|\] \[\left| \frac{-5}{5} \right|\] \[\left| -1\right|\] \[1\] So \[\left| \frac{7-12}{12-7} \right|=1\]

OpenStudy (anonymous):

woohoo!! Thank you!! I can check my odd probs, but can't check my even ones! Thnak you!